OpenStudy (anonymous):

Write the following system as a matrix equation: -3x+4y-3z=-20 -3x -3z=-12 x-2y-5z=-16

jimthompson5910 (jim_thompson5910):

-3x+4y-3z=-20 -3x -3z=-12 x-2y-5z=-16 as a matrix equation is \[\Large \begin{pmatrix} -3 & 4 & -3\\ -3 & 0 & -3\\ 1 & -2 & -5 \end{pmatrix} \begin{pmatrix} x\\ y\\ z \end{pmatrix} = \begin{pmatrix} -20\\ -12\\ -16 \end{pmatrix} \]

jimthompson5910 (jim_thompson5910):

Basically you write the coefficients in a matrix and you multiply it by the column vector of the 3 variables. That result will be equal to the column vector of the right hand side values.